FMARS-15 Sol Summary Report 24-07-2023

 In Sol Summary

Crew 15 Sol Summary Report 24-07-2023

Sol: 1

Summary Title: Preparing the Hab

Author’s name: D. Terry Trevino

Mission Status: All Nominal

Sol Activity Summary: After about 8 hours of sleep, our day began with a solid breakfast of cereal and crystalized eggs mixed with bacon. Every day, we awake knowing, we are a team of five and perhaps the only people on Devon Island. That precipitated a song from Commander Wheeler’s era, 1969, Good Morning Starshine, by Oliver.

Our crew commander, Andrew Wheeler, mapped our day out, cleaning the space suits, preparing for a practice EVA, and learning best practices for donning and doffing the analog space suits. Commander Wheeler discussed our objectives for the day.

Part of our daily routine will be starting up the generators, and we have lovingly named them Genny, (Yanmar) Bambi, (Yanmar), and Rachel (Generac), along with Colin (Coleman), and Thumper (Dead Yanmar) who is missing a tank.

Commander Wheeler led Chief Science Officer Drayson and XO/Mission Scientist Trevino down to the creek below the HMP landing strip to collect water for the next two days while the engineering team of Caleb Pool and Andy Greco worked on organizing their stations and Crew Engineer Pool continued to repair spacesuits.

Lunch included a rehydrated soup mixed in water used the previous day and canned “luncheon meat” gently browned with freshly baked bread.

We practiced meditation twice while monitoring our heart rate variability with our Garmin watches.

We are monitoring Hab CO2 on the upper level.

Look Ahead Plan:

Anomalies in work: A fuel line leak in Genny appeared during startup for evening meal prep, and apparently, the fuel line is fragile from 6 years of exposure to the arctic elements. During our water run today, Commander Wheeler noted that the water in the creek appeared much lower, and we are anticipating it will continue to run at a reduced rate.

Weather: Today was warm, bright, and sunny, with winds approaching .5 mph or 1 m/s. The temperature in the Hab hit 22C and outside 15C.

Crew Physical Status: Crew health is nominal, with no significant injuries other than the knicks and muscle strains from the strenuous work over the past nine days.

EVA: No EVAs have occurred to date.

Reports to be filed: Sol Summary, Operations Report, Journalist Report, EVA Report, EVA Request.

Support Requested: No support was requested other than previously mentioned in an email exchange between Crew Engineer Andy Greco and The Mars Society – Executive Director – James Burk

Start typing and press Enter to search